# Generative Ghosts: Anticipating Benefits and Risks of AI Afterlives

### v3

@@ 4. Benefits and Risks of Generative Ghosts
Having introduced the concept of generative ghosts and the key associated design parameters, we now discuss the potential benefits of thoughtfully and safely implemented generative ghosts, followed by a discussion of potential risks associated with various instantiations of this paradigm.
@@ 4.1. Generative Ghosts: Potential Benefits
We first consider likely motivations for creating generative ghosts. With appropriate attention to technical concerns (interaction design, AI safety, etc.) and sociotechnical concerns (ethics, cultural concerns, legal ramifications, etc.), one can envision a positive future in which generative ghosts offer a host of benefits both to the representee, the bereaved, and society more broadly.
@@ 4.1.1. Potential Benefits for Representees
Individuals may choose to create first-party generative ghosts as part of their end-of-life planning for several reasons. On a personal level, people may feel comforted by or even feel hopeful in the idea of a “digital afterlife.” Prior work on digital legacy has documented people’s motivations to capture their life stories (Thomas and Briggs, 2014), major life events (Jamison-Powell et al., 2016), and pass on values to future generations (Gulotta et al., 2017). Posthumous, personalized AI representations may likewise be seen as a way for people to ensure that they are remembered, whether by family, friends, or perhaps the world at large.
The likely nature of generative ghost architectures makes such remembrances higher-fidelity, more interactive, and more personalized than today’s more common digital memorials, such as memorialized social media accounts. Generative ghosts may help individuals feel agency over their posthumous future, such as by presenting an ability to comfort loved ones after they have passed, to offer advice to loved ones at future events (e.g., convey well wishes and advice upon the marriage of a great-grandchild), or to preserve their personal, religious, and/or cultural heritage by sharing knowledge and wisdom specific to their background and time in history to future generations.
@@ 4.1.2. Potential Benefits for the Bereaved
Interacting with a generative ghost may provide emotional support to close friends and family of the deceased. Prior research has considered the impact of online memorials on the bereaved (e.g. (Walter et al., 2012)), often responding to concerns that such spaces might prolong grief (see Section 4.2.1). Yet scholars cite the benefits of online memorials that allow loved ones to maintain healthy “continuing bonds” (Klass et al., 1996) with the deceased (e.g., (Getty et al., 2011; Brubaker and Hayes, 2011; Hjorth, 2021); c.f., (DeGroot, 2018)), often in a space where those grieving can gather (Carroll and Landry, 2010; Brubaker et al., 2013).
@@
One-sided conversations with deceased loved ones, such as speaking graveside or writing to the deceased on a memorial, are common and well-recognized in the literature (DeGroot, 2018; Brubaker and Hayes, 2011; Brubaker et al., 2013) and are generally considered a part of healthy grieving (Klass et al., 1996). As summarized by DeGroot (DeGroot, 2018), these conversations are cathartic and help the living confront their loss: “People essentially struggle to develop an identity where the deceased is now part of the past self and not the present self… and communicating with the deceased appears to ease the transition” (DeGroot, 2018). However, it is worth noting that the dead do not typically respond. Rather, people imagine how the deceased would reply (DeGroot, 2018). As such, it is unclear how responses from a generative ghost might impact bereavement.
@@ 4.1.3. Potential Benefits for Society
Generative ghosts may also be beneficial for smaller cultural groups. They may be one way to preserve the collective wisdom of elders, as well cultural heritage such as knowledge of dying languages, religions, or other small-membership cultural practices that are at risk of being forgotten as their members age. For instance, generative ghosts may be one way to preserve historical knowledge about events such as the Holocaust before the few remaining elderly survivors with firsthand experience pass away.
Indeed, the creation of generative ghosts may enrich the practice of disciplines such as museum curation, historical scholarship, anthropology, and other related humanities. These disciplines may benefit, not only from the static knowledge encoded in generative ghosts, but from the ability to interactively query such representations to understand perspectives and events in new ways. For example, devout adherents to a specific religious tradition might be able to converse with generative ghosts of deceased elders or scholars of their faith about changing aspects of society (e.g., perhaps an Orthodox Jew might engage with generative ghosts of Talmudic scholars to debate the pros and cons of whether kashrut allows consumption of synthetic, lab-grown pork). Generative ghosts might also represent archetypes or amalgamations rather than specific individuals, particularly when developed from historical records (e.g., “citizen of Pompeii,” “typical resident of Colonial Williamsburg”).
@@ 4.2.1. Mental Health Risks
Earlier we discussed ways in which generative ghosts might offer mental health benefits to mourners. However, ghosts may also pose mental health risks for the bereaved. Scholars in thanatology and clinical psychology have long distinguished between adaptive coping – strategies that integrate the loss into an individual’s life, fostering emotional adjustment, well-being, and personal growth over time – and maladaptive coping – behavior that may obstruct healthy grieving, prolonging distress and resulting in depression, anxiety, and complicated grief (also known as Prolonged Grief Disorder)(Stroebe et al., 2013; Shear, 2015). Meanwhile, studies of prior generations of AI technologies (and even of non-AI computing systems (Reeves and Nass, 1996)) have found that people tend to anthropomorphize such systems and even develop attachments (Gabriel et al., 2024). As such, ghosts may include risks related to delayed accommodation, excessive attachment, information overload, anthropomorphism, deification, and second deaths. We anticipate that without careful design generative ghosts might magnify these risks due to the increasing fidelity of generative AI technologies and the modeling of the agent on a specific individual.
@@
A related concern is the possibility of developing excessive attachment to generative ghosts, as indicated by spending a disproportionate amount of time interacting with representations of the deceased in lieu of interacting with the living. Already, there are indications that some individuals find AI-powered companions, such as those from Replika (replika.ai), to be highly compelling (Liang, 2023); generative ghosts’ basis in real, beloved individuals rather than in fictional personas or historical figures may amplify this type of attachment.
@@
In addition to spending an unhealthy amount of time interacting with a generative ghost, another aspect of excessive attachment may take the form of consulting a generative ghost’s advice for momentous or even mundane life decisions. While soliciting input from ghosts might be positive or harmless in some circumstances, the availability of ghosts might also result in information overload, lead to choice paralysis, or create tensions over how to weight competing recommendations from a generative ghost vs. living friends and relatives.
@@
Anthropomorphism of generative ghosts is a potential risk for mourners, if they become convinced that a generative ghost truly is the deceased (rather than a computer program representing them). Anthropomorphism might increase the likelihood of other mental health risks such as excessive attachment or delayed accommodation. Some design choices we discussed earlier may increase anthropomorphism risk, particularly the use of reincarnation metaphors, evolving representations, and physical embodiments.
@@
Deification is a more extreme version of anthropomorphism risk, in which a mourner might develop religious or supernatural beliefs about a generative ghost. Such beliefs may carry risks such as alienating the mourner from living companions, excessive attachment, over-reliance on the generative ghost for life advice (i.e., treating it as an oracle), and/or a propensity to unquestioningly carry out actions suggested by the generative ghost (potentially including actions detrimental to themselves or others).
@@
An additional mental health risk is that of second deaths or second losses, the experience of grieving a loved one again. [^1] Grief scholarship sometimes uses “secondary loss” to refer to experiences of grief that follow the “primary loss” of a loved one (e.g., loss of relationships, home, shared activities). Here, however, we focus on circumstances that might result in people engaging with their primary loss and grief a second time. In digital contexts, second deaths occur when data becomes unavailable either through technical obsolescence, deletion, or lack of access (Bassett, 2022). In the case of generative ghosts, second death refers to the emotional harm that may result if a ghost ceases to exist. Second deaths may occur for myriad reasons, including economic (i.e., the company that maintains the generative ghost service goes out of business; survivors’ or the deceased’s estate can no longer afford maintenance fees), regulatory (i.e., a government outlaws generative ghosts, which may be particularly likely in theocracies where AI afterlives threaten established belief systems), technological (i.e., obsolete generative ghosts no longer run on future infrastructure), or security reasons (i.e., a hacker disables a generative ghost). While second loss is not unique to AI systems, the potential intensity of engagement that generative ghosts can support means that the loss experienced at the disabling of a generative ghost might be especially profound.
@@ 4.2.2. Reputational Risks
Privacy Risks refer to scenarios in which a generative ghost exposes true information that the deceased would not have wanted to be revealed (e.g., because the information is embarrassing, criminal, hurtful to loved ones, etc.). While some privacy risks may be anticipated and prevented via rules when creating a ghost (i.e., “don’t tell my spouse about my affair”), revelations that arise from generated content may be more difficult to prevent (e.g., if the AI correctly infers and reveals the deceased’s sexual orientation based on patterns in their data, even though they were closeted while alive). Likewise, what constitutes a privacy risk changes over time. Aspects of the deceased’s life or beliefs that were socially acceptable during their lifetime may come to be frowned upon by future societies (e.g., perhaps in an environmentally conscious future, people who consumed certain kinds of meat or used certain levels of fossil fuels might be retroactively viewed as immoral).
@@ 4.2.4. Socio-cultural Risks
Finally, we note that generative ghosts may introduce socio-cultural risks. While extremely important, this category of risk is also particularly speculative as it is difficult to anticipate how novel technologies may impact society at scale. For example, widespread adoption of generative ghosts might cause profound changes to cornerstones of modern society such as the labor market, interpersonal relationships, or religious institutions. While it is difficult to anticipate such societal impacts, even recognizing the possibility of systemic societal change without being confident of its precise direction can help us to be more thoughtful in the design of novel technologies.
@@
Another example of how generative ghosts may impact socio-cultural practices is in how they may alter interpersonal relationships and attendant social structures. For example, in the past two decades we have seen profound changes to social practices resulting from the widespread adoption of novel technologies such as smartphones and social media. The age of AI agents is likely to result in similarly disruptions of status quo socialization and relationship patterns (Shavit et al., 2023), and generative ghosts may be more disruptive to existing relationship patterns than general-purpose agents because they are modeled on specific loved ones rather than fictional characters or archetypes, possibly making them more compelling and potentially more likely to substitute for traditional relationships.
@@ 5. Discussion
In this paper, we defined the concept of generative ghosts, agentic griefbots that may execute a variety of behaviors and actions on behalf of a deceased individual, including a detailed taxonomy of the many design dimensions such technologies may entail. Although digital memorials and simple griefbots have existed for years, generative AI is fundamentally changing the nature and scale of postmortem representations, making generative ghosts a distinct and novel source of AI ethics concerns and a likely cause of social change. As documented in the Related Work section, the recent availability of powerful generative AI models has led to an increase in tech-savvy individuals creating generative-ghost-like systems and startup companies purporting to offer such services; the accelerating interest in post-mortem agents suggests this is an important area for technologists, ethicists, and policy makers to anticipate, monitor, and address. We contributed a taxonomy of key design dimensions of generative ghosts, and used this taxonomy to systematically identify potential benefits and risks of this new paradigm of post-mortem representation. Here, we reflect on the importance of careful partnership between the AI and HCI communities to design generative ghosts in a manner that mitigates risk, discuss the need to develop policies that protect values around privacy, consent, and safety, discuss the complexities of foreseeing the societal impacts of “AI Afterlives,” and identify key avenues for future research.
@@ 5.1. Interfaces to Mitigate Risk
Earlier in this paper we introduced several design dimensions for generative ghosts. Choices regarding these design dimensions may influence the benefit/risk landscape of this emerging technological paradigm.
@@
Careful attention to the interaction design and interfaces of generative ghosts is vital for ensuring that these systems empower representees and the bereaved, and maximize the likelihood of socially beneficial outcomes over risky ones. This includes investing in user studies and social science research to more carefully understand aspects of the design space, such as what interfaces and interactions increase anthropomorphism risks (Abercrombie et al., 2023; Shanahan, 2023; Gabriel et al., 2024; McIlroy-Young et al., 2022; Lee et al., 2023), and what factors (i.e., features of generative ghosts, contexts of loss, attributes of the bereaved) may contribute to mental health risks from engaging with such technologies.
@@
Whether a ghost is designed to act as a “reincarnation” of or “represents” the deceased is a particularly important feature of the design space to consider. Designers will have to make choices about whether a ghost speaks as the representee, assuming their voice (which some prior work (Brubaker et al., 2013) and case studies (Newton, 2016) suggest can be unsettling) vs. speaking about the representee (ostensibly from the point of view of the system or bot). Embodiment can present similar issues, particularly if the ghost’s physical design resembles the representee.
@@
Yet we do not mean to suggest that there are clear right and wrong approaches for each feature in our design space. After all, a ghost that reincarnates a spouse is likely experienced quite differently than one that reincarnates a historical and/or public figure such as Abraham Lincoln. We suspect that reincarnation-style or embodied interfaces for historical figures carry less risk than for personal remembrances. Characteristics of the representee, particularly in relationship to the user, will ultimately have profound impacts on how users experience ghosts. However, interaction design can play an important role in framing these experiences.
@@
Given the sensitive nature of generative ghosts, it is particularly important to be vigilant against the use of dark patterns (Narayanan et al., 2020) in the design of these technologies, to minimize the risk of, for instance, the formation of addictive parasocial relationships that might harm mental health. For instance, one might consider what the equivalent of undesirable “push notifications” might be for an agentic system such as a generative ghost – perhaps ghosts should only respond to interactions initiated by the living rather than initiating interactions.
@@
Generative ghosts might even include interface designs to proactively guard against likely harms. For instance, ghosts might monitor the patterns of interaction with the bereaved and analyze these interactions for patterns of overuse or other types of harm; in such cases, the systems might initiate interventions such as offering referrals to mental health professionals, reducing its own fidelity, reducing the hours during which it is available, etc. The types of use patterns that might prompt concern and the appropriate system responses are important areas for future study.

### v4

@@ 4. Benefits and Risks of Generative Ghosts
Having introduced the concept of generative ghosts and the dimensions associated with this design space, we now discuss the potential benefits of thoughtfully and safely implemented generative ghosts, followed by a discussion of potential risks associated with various implementations.
@@ 4.1. Generative Ghosts: Potential Benefits
We first consider likely motivations for creating generative ghosts. With appropriate attention to technical concerns (e.g., interaction design, AI safety, etc.) and sociotechnical concerns (e.g., ethics, cultural concerns, legal ramifications, etc.), one can envision a positive future in which generative ghosts offer a host of benefits both to the representee, the bereaved, and society more broadly.
@@ 4.1.1. Potential Benefits for Representees
Individuals may choose to create first-party generative ghosts as part of their end-of-life planning for several reasons. On a personal level, people may feel comforted by or even feel hopeful about the idea of a “digital afterlife.” Prior work on digital legacy has documented people’s motivations to capture their life stories (Thomas and Briggs, 2014), major life events (Jamison-Powell et al., 2016), and pass on values to future generations (Gulotta et al., 2017). Posthumous, personalized AI representations may likewise be seen as a way for people to ensure that they are remembered, whether by family, friends, or perhaps the world at large.
The likely nature of generative ghost architectures makes such remembrances higher-fidelity, more interactive, and more personalized than today’s more common digital memorials, such as memorialized social media accounts. As such, generative ghosts may help give individuals a sense of agency over their posthumous future, such as by presenting an ability to comfort loved ones after they have passed, to offer advice to loved ones at future events (e.g., convey well wishes and advice upon the marriage of a great-grandchild), or to preserve their personal, religious, and/or cultural heritage by sharing knowledge and wisdom specific to their background and time in history to future generations.
@@ 4.1.2. Potential Benefits for the Bereaved
Interacting with a generative ghost may provide emotional support to close friends and family of the deceased. Prior research has considered the impact of online memorials on the bereaved (e.g. (Walter et al., 2012)), often responding to concerns that such spaces might prolong grief (see Section 4.2.1). Yet scholars cite the benefits of online memorials that allow loved ones to maintain important “continuing bonds” (Klass et al., 1996) with the deceased (e.g., (Getty et al., 2011; Brubaker and Hayes, 2011; Hjorth, 2021); c.f., (DeGroot, 2018)), often in a space where those grieving can gather (Carroll and Landry, 2010; Brubaker et al., 2013).
@@
One-sided conversations with deceased loved ones, such as speaking graveside or writing to the deceased on a memorial, are common and well-recognized in the literature (DeGroot, 2018; Brubaker and Hayes, 2011; Brubaker et al., 2013) and are generally considered a part of healthy grieving across many grief paradigms (e.g., (Klass et al., 1996)). As summarized by DeGroot (DeGroot, 2018), these conversations are cathartic and help the living confront their loss: “People essentially struggle to develop an identity where the deceased is now part of the past self and not the present self… and communicating with the deceased appears to ease the transition” (DeGroot, 2018). However, it is worth noting that the dead do not typically respond. Rather, people imagine how the deceased would reply (DeGroot, 2018). As such, it is unclear how a generative ghost that can respond might impact bereavement.
@@ 4.1.3. Potential Benefits for Society
Generative ghosts may also be beneficial for smaller cultural groups. They may be one way to preserve the collective wisdom of elders, as well as cultural heritage including the knowledge of dying languages, religions, or other small-membership cultural practices that are at risk of being forgotten as their members age. For instance, generative ghosts may be one way to preserve historical knowledge about events such as the Holocaust before the few remaining elderly survivors with firsthand experience pass away (Zucker et al., [n. d.]; Colavizza et al., [n. d.]).
Indeed, the creation of generative ghosts may enrich the practice of disciplines such as museum curation, historical scholarship, anthropology, and other related humanities. These disciplines may benefit, not only from the static knowledge encoded in generative ghosts, but from the ability to interactively query such representations to understand perspectives and events in new ways. For example, devout adherents to a specific religious tradition might be able to converse with generative ghosts of deceased elders or scholars of their faith about changing aspects of society (e.g., perhaps an Orthodox Jew might engage with generative ghosts of Talmudic scholars to debate the pros and cons of whether kashrut allows consumption of synthetic, lab-grown pork). Generative ghosts might also represent archetypes or amalgamations rather than specific individuals, particularly when developed from historical records (e.g., “citizen of Pompeii,” “typical resident of Colonial Williamsburg,” etc.).
@@ 4.2.1. Mental Health Risks
Earlier we discussed ways in which generative ghosts might offer mental health benefits to mourners. However, ghosts may also pose mental health risks for the bereaved, some of which we outline in this section. It is important to note, however, that what is considered a mental health risk is context-dependent as grief paradigms vary extensively across cultures. For example, common Chinese rituals and practices related to ancestors — practices that appear to be shaping the acceptability of generative ghosts already — might be interpreted as problematic elsewhere. Similarly, the Māori of New Zealand see ancestors as kaitiaki (guardians) who offer spiritual support and ensure the well-being of their descendants (Schwass and of New Zealand, [n. d.]). As such, behavior considered risky in a Western context may be seen as normal or even beneficial in another. The mental health risks we outline in this section are predominantly drawn from Western perspectives as they reflect the concerns most frequently raised in the literature related to technology and mortality. However, even as we enumerate potential risks, we would encourage readers to consider cultural factors that shape what is considered a risk vs. a benefit.
Western scholars in thanatology and clinical psychology have long distinguished between adaptive coping – strategies that integrate the loss into an individual’s life, fostering emotional adjustment, well-being, and personal growth over time – and maladaptive coping – behavior that may obstruct healthy grieving, prolonging distress and resulting in depression, anxiety, and complicated grief (also known clinically as Prolonged Grief Disorder)(Stroebe et al., 2013; Shear, 2015). Meanwhile, studies of prior generations of AI technologies (and even of non-AI computing systems (Reeves and Nass, 1996)) have found that people tend to anthropomorphize such systems and even develop attachments (Gabriel et al., 2024). As such, ghosts may include risks related to delayed accommodation, complicated grief, information overload, anthropomorphism, deification, and second deaths, which we detail below. We anticipate that without careful design that generative ghosts might magnify risks due to the increasing fidelity of generative AI technologies and the modeling of the agent on a specific individual.
@@
A related concern is the possibility of complicated grief, i.e., developing an overly strong reliance on generative ghosts as indicated by spending a disproportionate amount of time engaging with representations of the deceased rather than interacting with the living. Already, there are indications that some individuals find AI-powered companions, such as those from Replika (replika.ai), to be highly compelling (Liang, 2023); generative ghosts’ basis in real, beloved individuals rather than in fictional personas or historical figures may amplify addictive risks of interacting with AI agents (Mahariarchive and Pataranutaporn, 2024).
@@
People may also overly rely on a generative ghost’s advice for momentous or even mundane life decisions. While soliciting input from ghosts might be positive or harmless in some circumstances, one can imagine how the use of ghosts could result in information overload, lead to choice paralysis, or create tensions over how to weigh competing recommendations from a generative ghost vs. living friends and relatives.
@@
Anthropomorphism of generative ghosts is a potential risk for mourners, if they become convinced that a generative ghost truly is the deceased (rather than a computer program representing them). Anthropomorphism might increase the likelihood of other mental health risks such as complicated grief or delayed accommodation (Shear, 2015). Some design choices we discussed earlier may increase anthropomorphism risk, particularly the use of reincarnation metaphors, evolving representations, and physical embodiments.
@@
Deification is a more extreme version of the anthropomorphism risk, in which a mourner might develop religious or supernatural beliefs about a generative ghost that are culturally atypical. Such beliefs may carry risks such as alienating the mourner from living companions, delayed accommodation, over-reliance on the generative ghost for life advice (i.e., treating it as an oracle), and/or a propensity to unquestioningly carry out actions suggested by the generative ghost (potentially including actions detrimental to themselves or others).
@@
An additional mental health risk is that of second deaths or second losses, the experience of grieving a loved one again. [^8] Grief scholarship sometimes uses “secondary loss” to refer to experiences of grief that follow the “primary loss” of a loved one (e.g., loss of relationships, home, shared activities). Here, however, we focus on circumstances that might result in people engaging with their primary loss and grief a second time. In digital contexts, second deaths occur when data becomes unavailable either through technical obsolescence, deletion, or lack of access (Bassett, 2022), such as the sudden deprecation of MySpace features that effectively eliminated memorial messages from profiles (Brubaker, 2015). In the case of generative ghosts, second death refers to the emotional harm that may result if a ghost ceases to exist. Second deaths may occur for myriad reasons, including economic (i.e., the company that maintains the generative ghost service goes out of business; survivors’ or the deceased’s estate can no longer afford maintenance fees), regulatory (i.e., a government outlaws generative ghosts, which may be particularly likely in theocracies where AI afterlives threaten established belief systems), technological (i.e., obsolete generative ghosts no longer run on future infrastructure), or security reasons (i.e., a hacker disables a generative ghost). While second loss is not unique to AI systems, the potential intensity of engagement that generative ghosts can support means that the loss experienced at the disabling of a generative ghost might be especially profound.
@@ 4.2.2. Reputational Risks
Privacy Risks refer to scenarios where a generative ghost exposes true information that the deceased would not have wanted to be revealed (e.g., because the information is embarrassing, criminal, hurtful to loved ones, etc.). While some privacy risks may be anticipated and prevented via rules when creating a ghost (i.e., “don’t tell my spouse about my affair”), revelations that arise from generated content may be more difficult to prevent (e.g., if the AI correctly infers and reveals the deceased’s sexual orientation based on patterns in their data, even though they were closeted while alive). Likewise, what constitutes a privacy risk changes over time. Aspects of the deceased’s life or beliefs that were socially acceptable during their lifetime may come to be frowned upon by future societies (e.g., perhaps in an environmentally conscious future, people who consumed certain kinds of meat or used certain levels of fossil fuels might be retroactively viewed as immoral).
@@ 4.2.4. Socio-cultural Risks
Finally, we note that generative ghosts may introduce socio-cultural risks. While extremely important, this category of risk is also particularly speculative as it is difficult to anticipate how novel technologies may impact society at scale. For example, the widespread adoption of generative ghosts might cause profound changes to cornerstones of modern society such as the labor market, interpersonal relationships, or religious institutions. While it is difficult to anticipate such societal impacts, even recognizing the possibility of systemic societal change without being confident of its precise direction can help us to be more thoughtful in the design of novel technologies.
@@
Another example of how generative ghosts may impact socio-cultural practices is in how they may alter interpersonal relationships and attendant social structures. For example, in the past two decades, we have seen profound changes in social practices resulting from the widespread adoption of novel technologies such as smartphones and social media. The age of AI agents is likely to result in similar disruptions of status quo socialization and relationship patterns (Shavit et al., 2023), and generative ghosts may be more disruptive to existing relationship patterns than general-purpose agents because they are modeled on specific loved ones rather than fictional characters or archetypes, possibly making them more compelling and potentially more likely to substitute for traditional relationships.
@@ 5. Discussion
In this paper, we defined the concept of generative ghosts, agentic griefbots that may execute a variety of behaviors and actions on behalf of a deceased individual, including a design space of the many design dimensions such technologies may entail. Although digital memorials and simple griefbots have existed for years, generative AI is fundamentally changing the nature and scale of postmortem representations, making generative ghosts a distinct and novel source of AI ethics concerns and a likely cause of social change. As documented in the Related Work section, the recent availability of powerful generative AI models has led to an increase in tech-savvy individuals creating generative-ghost-like systems and startup companies purporting to offer such services; the accelerating interest in post-mortem agents suggests this is an important area for technologists, ethicists, and policymakers to anticipate, monitor, and address. We contributed a design space of key design dimensions of generative ghosts and used this analytic framework to systematically identify potential benefits and risks of this new paradigm of post-mortem representation. Here, we reflect on the importance of careful partnership between the AI and HCI communities to design generative ghosts in a manner that mitigates risk, discuss the need to develop policies that protect values around privacy, consent, and safety, discuss the complexities of foreseeing the societal impacts of “AI Afterlives,” and identify key avenues for future research.
@@ 5.1. Interfaces to Mitigate Risk
Choices regarding the design dimensions we outline in this paper may influence the benefit/risk landscape of this emerging design space. Careful attention to the interaction design and interfaces of generative ghosts is vital for ensuring that these systems empower representees and the bereaved, and maximize the likelihood of socially beneficial outcomes over risky ones. This includes investing in user studies and social science research to more carefully understand aspects of the design space, such as what interfaces and interactions increase anthropomorphism risks (Abercrombie et al., 2023; Shanahan, 2023; Gabriel et al., 2024; McIlroy-Young et al., 2022; Lee et al., 2023), and what factors (i.e., features of generative ghosts, contexts of loss, attributes of the bereaved) may contribute to mental health risks from engaging with such technologies.
@@
Whether a ghost is designed to act as a “reincarnation” of or “represents” the deceased is a particularly important feature of the design space to consider. Designers will have to make choices about whether a ghost speaks as the representee, assuming their voice (which some prior work (Brubaker et al., 2013) and case studies (Newton, 2016) suggest can be unsettling to some) vs. speaking about the representee (ostensibly from the point of view of the system or bot). Embodiment can present similar issues, particularly if the ghost’s physical design resembles the representee.
@@
Yet we do not mean to suggest that there are clear right and wrong approaches for each dimension of our design space. After all, a ghost that reincarnates a spouse is likely experienced quite differently than one that reincarnates a historical and/or public figure such as Abraham Lincoln. We suspect that reincarnation-style and embodied interfaces for historical figures carry less risk than for personal friends and family. Characteristics of the representee, the user, and the relationship between the two, will ultimately have profound impacts on how users experience ghosts. However, interaction design can play an important role in framing these experiences.
@@
Given the sensitive nature of generative ghosts, it is particularly important to be vigilant against the use of dark patterns (Narayanan et al., 2020) in the design of these technologies, to minimize the risk of, for instance, the formation of addictive parasocial relationships that might harm mental health. For instance, one might consider what the equivalent of undesirable “push notifications” might be for a generative ghost. Perhaps ghosts should only respond to interactions initiated by the living rather than initiating interactions.
@@
Generative ghosts might even include interface designs to proactively guard against likely harms. For instance, ghosts might monitor the patterns of interaction with the bereaved and analyze these interactions for patterns of overuse or other types of harm. One could imagine scenarios where a system might initiate interventions such as offering referrals to mental health professionals, reducing its own fidelity, or reducing the hours during which it is available. The types of usage patterns that might prompt concern and the appropriate system responses are important areas for future study.
Another consideration for interface and interaction design is around transparency (i.e., whether and how to clearly signal to end-users that the agent is representing someone who is deceased rather than someone who is still living (Brubaker and Callison-Burch, 2016)). Concerns about transparency are already surfacing in adjacent spaces. For example, in 2024 a Polish Radio station broadcasted an interview with Nobel laureate Wislawa Szymborska without revealing that it was a simulated interview with an AI representation of Szymborska, who had died more than a decade earlier (Higgens, 2024). In China, a company called Super Brain provides a service in which AI replicas of deceased people can make phone or video calls to family members. According to one report, some customers use this service to avoid having to tell an older relative the distressing news that someone has died (Yang, 2024b). Perhaps AI agents based on real people (e.g., ”generative clones”) should utilize digital watermarking or fingerprinting capabilities that enable interaction partners to verify whether a representee is alive or deceased. Similar technology might also be used to verify whether an agent is or is not authorized by the representee.